1 angle is 60 another 20 whats the missing angle

Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

If this is a triangle, then the missing angle is 180 - 60 - 20 = 100 degrees

Answer by stanbon(75887)

1 angle is 60 another 20 whats the missing angle
------
Assuming you are referring to the angles of a triangle,
the sum of the angles must be 180 degrees.
-------
Equation:
60 + 20 + x = 180
x + 80 = 180
-----
x = 100 degrees (the 3rd angle)
=============================
Cheers,
Stan H.
===============